# Clarifying the effect of mean subtraction on Dynamic Mode Decomposition

### v5

@@ D.2 Equivalence when $\mathbf{X}_{\rm ms}$ has linearly dependent columns
Once again, we can use Eqs. 46 and 45 to introduce $\boldsymbol{\mathbf{l}}^{\perp}$ and leverage its orthogonality to the first $n-1$ columns of $\mathbf{\Theta}_{1}$:
$\underbrace{\tilde{\mathbf{C}}_{0}(\mathbf{\Lambda}-\mathbf{I})}\mathbf{\Theta}_{1}\begin{bmatrix}n\\ n-1\\ \vdots\\ 1\end{bmatrix}\left(\frac{-1}{n+1}\right)~{}={}$ $\underbrace{(\boldsymbol{\mathbf{l}}^{\perp})^{H}\mathbf{\Theta}_{1}\begin{bmatrix}n\\ n-1\\ \vdots\\ 1\end{bmatrix}}\left(\frac{-1}{n+1}\right)$
@@
Since we’ve already seen that $\boldsymbol{\mathbf{l}}^{\perp}\neq\boldsymbol{\mathbf{0}}$, the matrix-vector product $\mathbf{X}_{\rm ms}\boldsymbol{\mathbf{e}}_{1}\neq\boldsymbol{\mathbf{0}}$ and, consequently, we have DMD-DFT equivalence.
D.3 A modal decomposition of mean-subtracted data
When the Koopman mode expansion possesses only a finite number of terms, Lemma A.1 tells us that the time series $\mathbf{Z}$ can be written as the product of $\mathbf{C}$ and $\mathbf{\Theta}$. Here, we derive an analogous factorization for the mean-removed data-set $\mathbf{Z}_{\rm ms}$.
Definition D.7 (A conformal partitioning of $\mathbf{C}$ and $\mathbf{\Theta}$).
Suppose (3.8) and (3.10) are true. Additionally, let $1$ be a Koopman eigenvalue i.e., $1\in\{\,\lambda_{i}\,\}_{i=1}^{r}.$ Then, Eq. 24 motivates defining $\boldsymbol{\mathbf{c}}_{1}$, $\mathbf{C}_{\not\owns 1}$ and $\mathbf{\Theta}_{\not\owns 1}$ via the following conformal partitioning:
@@ D.3 A modal decomposition of mean-subtracted data
Table 3: Constructing $\mathbf{C}_{\rm ms}$ and $\mathbf{\Theta}_{\rm ms}$: There are four possible scenarios in mean-subtraction, depending on the presence of a Koopman eigenvalue at 1 (Branch 1) and the value of $\boldsymbol{\mathbf{\mu}}$ (Branch 2). The peculiarities of each case are incorporated in the definitions of $\mathbf{C}_{\rm ms}$ and $\mathbf{\Theta}_{\rm ms}$ (last two columns), with an eye towards inducing an analogue of Eq. 24 for $\mathbf{Z}_{\rm ms}$.
Case Branch 1 Branch 2 $\mathbf{C}_{\rm ms}$ $\mathbf{\Theta}_{\rm ms}$
I $1\notin\{\,\lambda_{i}\,\}_{i=1}^{r}$ $\boldsymbol{\mathbf{\mu}}=\boldsymbol{\mathbf{0}}$ $\mathbf{C}$ $\mathbf{\Theta}$
II ′′ $\boldsymbol{\mathbf{\mu}}\neq\boldsymbol{\mathbf{0}}$ $\begin{bmatrix}\mathbf{C}&-\boldsymbol{\mathbf{\mu}}\end{bmatrix}$ $\begin{bmatrix}\mathbf{\Theta}\\ \mathbf{1}_{n+1}^{T}\end{bmatrix}$
III $1\in\{\,\lambda_{i}\,\}_{i=1}^{r}$ $\boldsymbol{\mathbf{\mu}}=\boldsymbol{\mathbf{c}}_{1}$ $\mathbf{C}_{\not\owns 1}$ $\mathbf{\Theta}_{\not\owns 1}$
IV ′′ $\boldsymbol{\mathbf{\mu}}\neq\boldsymbol{\mathbf{c}}_{1}$ $\begin{bmatrix}\boldsymbol{\mathbf{c}}_{1}-\boldsymbol{\mathbf{\mu}}&\mathbf{C}_{\not\owns 1}\end{bmatrix}$ $\begin{bmatrix}\boldsymbol{\mathbf{1}}_{n+1}^{T}\\ \mathbf{\Theta}_{\not\owns 1}\end{bmatrix}$
Lemma D.8.
Suppose (3.8) and (3.10) are true. If we construct the two matrices $\mathbf{C}_{\rm ms}$ and $\mathbf{\Theta}_{\rm ms}$ as described by Definition D.7 and Table 3, then, every column of $\mathbf{C}_{\rm ms}$ is non-zero, $\mathbf{\Theta}_{\rm ms}$ has distinct nodes and, more importantly,
@@
Proof D.9.
A direct application of Lemma A.1 to Eq. 10, in the context of Definition D.7 and Table 3.
Remark D.10.
@@
D.4 Over-sampling and linear consistency prevent equivalence
@@
Table 4: Delineating the effect of mean subtraction: There are four possible outcomes of mean-removal, depending on the presence (or absence) of $1$ in the nodes of $\mathbf{\Theta}_{\rm ms}$ (Branch 1) and in the Koopman eigen-values (Branch 2). For each case, we detail $\sigma(\mathbf{\Theta}_{\rm ms})$ alongside $\mathbf{C}_{\rm ms}$ and $\mathbf{\Theta}_{\rm ms}$ (whenever they have a simple expression).
Case Branch 1 Branch 2 $\sigma(\mathbf{\Theta}_{\rm ms})$ $\mathbf{C}_{\rm ms}$ $\mathbf{\Theta}_{\rm ms}$
@@ II $1\in\{\,\lambda_{i}\,\}_{i=1}^{r}$$\{\,\lambda_{i}\,\}_{i=1}^{r}$ $\mathbf{\Theta}$
Firstly, we show that a Koopman invariant dictionary ensures a non-zero mean $\boldsymbol{\mathbf{\mu}}$, except in Case III. This is proven in three parts and begins by studying Cases II and IV:
Proposition D.12.
Suppose (3.8) and (3.10) are true. If $\tilde{\mathbf{C}}$ has full column rank and $1$ is a Koopman eigenvalue i.e., $1\in\{\,\lambda_{i}\,\}_{i=1}^{r}$, then, the temporal mean $\boldsymbol{\mathbf{\mu}}\neq\boldsymbol{\mathbf{0}}.$
Proof D.13.
When (3.8) is true, the time-average $\boldsymbol{\mathbf{\mu}}$, which is defined in Eq. 9, can be expanded using Eq. 24 as follows:
@@
Now, $1\in\{\,\lambda_{i}\,\}_{i=1}^{r}$ ensures that $\mathbf{\Theta}\mathbf{1}_{n+1}$ has at least one non-zero coefficient and, hence, is non-zero. Furthermore, (3.10) lets the full column rank of $\tilde{\mathbf{C}}$ be inherited by $\mathbf{C}$. Combining these two inferences, we get $\boldsymbol{\mathbf{\mu}}\neq\boldsymbol{\mathbf{0}}$.
Case I is tackled by finding a necessary condition for zero mean and taking the contrapositive.
Proposition D.14.
Suppose (3.8) and (3.10) are true. If the temporal mean $\boldsymbol{\mathbf{\mu}}=\boldsymbol{\mathbf{0}}$ and $1\in\sigma(\mathbf{\Theta}_{\rm ms})$, then $\tilde{\mathbf{C}}$ has linearly dependent columns.
Proof D.15.
By (3.10), it is sufficient to establish that the vector,
$\boldsymbol{\mathbf{z}}:=\mathbf{\Theta}\mathbf{1}_{n+1},$
is a non-trivial element of $\mathcal{N}(\mathbf{C})$.
Firstly, by the definitions of $\mathbf{C}_{\rm ms}$ and $\mathbf{\Theta}_{\rm ms}$ in Table 3, we have,
$\boldsymbol{\mathbf{\mu}}~{}=~{}\boldsymbol{\mathbf{0}}~{}\implies~{}\mathbf{C}_{\rm ms}=\mathbf{C},~{}\mathbf{\Theta}_{\rm ms}=\mathbf{\Theta}.$
As a result, the condition $1\in\sigma(\mathbf{\Theta}_{\rm ms})$ translates, by Eq. 22, to $1\in\{\,\lambda_{i}\,\}_{i=1}^{r}$, which once again gives $\mathbf{\Theta}\mathbf{1}_{n+1}=\boldsymbol{\mathbf{z}}\neq\boldsymbol{\mathbf{0}}$. Consequently, unpacking the action of $\mathbf{C}$ on $\boldsymbol{\mathbf{z}}$ yields the desired conclusion:
$\mathbf{C}\underbrace{\boldsymbol{\mathbf{z}}}\frac{1}{n+1}~{}=~{}\underbrace{\mathbf{C}\mathbf{\Theta}\mathbf{1}_{n+1}\frac{1}{n+1}}~{}=~{}\boldsymbol{\mathbf{\mu}}~{}=~{}\boldsymbol{\mathbf{0}}\implies\boldsymbol{\mathbf{z}}\in\mathcal{N}(\mathbf{C}).$
Compiling the above results and studying Case III, we find that it is the only scenario when the mean is $\boldsymbol{\mathbf{0}}$.
Lemma D.16.
Suppose (3.8) and (3.10) are true. If $\tilde{\mathbf{C}}$ has full column rank, then, the mean $\boldsymbol{\mathbf{\mu}}$ is zero only in Case III of Table 4.
Proof D.17.
We prove the forward implication by contraposition. Consider Cases I, II and IV in Table 4. By Proposition D.12, the mean is non-zero for Cases II and IV. In Case I, we have $1\in\sigma(\mathbf{\Theta}_{\rm ms})$. Since $\tilde{\mathbf{C}}$ has linearly independent columns, we can apply the contrapositive of Proposition D.14 to show $\boldsymbol{\mathbf{\mu}}\neq\boldsymbol{\mathbf{0}}$.
The backward implication is proven by contradiction. Suppose we have Case III and $\boldsymbol{\mathbf{\mu}}\neq\boldsymbol{\mathbf{0}}$. Since $1\notin\{\,\lambda_{i}\,\}_{i=1}^{r}$, removing the non-zero mean from $\mathbf{Z}$ will simply cause $1\in\sigma(\mathbf{\Theta}_{\rm ms})$ which is in contradiction with Case III.
@@
Proposition D.18.
Suppose (3.8) and (3.10) are true. Then,
@@
Technically, we are now embarking on the proof of Theorem 4.5. However, the arguments involved are also useful in establishing Corollary 4.8. So, they have been organized into the two theorems below to facilitate their reuse. Both of these are proven by contradiction and we begin by studying Cases III and IV.
Theorem D.20.
Suppose (3.8) and (3.10) are true. Additionally, assume that $1\notin\sigma(\mathbf{\Theta}_{\rm ms})$. If the training data is well-sampled i.e., $n\geq r$, then mean-subtracted DMD is not DFT for Case IV in Table 4. Under the stronger condition of over-sampling $(n\geq r+1)$, non-equivalence holds for Case III too.
Proof D.21.
We prove both cases by contradiction. Suppose mean-subtracted DMD is equivalent to DFT, as detailed in Definition 3.7. Then, by Theorem 4.1, we have:
@@
Hence, $\mathbf{\Theta}_{\rm aug}:=\begin{bmatrix}(\mathbf{\Theta}_{\rm ms})_{1}\\ \mathbf{1}_{n}^{H}\end{bmatrix}$ has linearly dependent rows.
But, the lower bounds on $n$ mean that $\mathbf{\Theta}_{\rm aug}$ has full row rank. Firstly, $1\notin\sigma(\mathbf{\Theta}_{\rm ms})$ ensures that $\mathbf{\Theta}_{\rm aug}$ has distinct nodes. While it has $n$ columns, the number of rows is case dependent. In Case III, $\mathbf{\Theta}_{\rm aug}$ has $r+1$ rows and at least $r+1$ columns. Similarly, in Case IV, $\mathbf{\Theta}_{\rm aug}$ has $r$ rows and at least $r$ columns. Hence, in both cases, Eq. 32 tells us that $\mathbf{\Theta}_{\rm aug}$ has full row rank which is a contradiction.
Cases I and II have a similar proof strategy where Proposition D.18 plays a crucial role.
Theorem D.22.
Suppose (3.8) and (3.10) are true. Additionally, assume that we have $1\in\sigma(\mathbf{\Theta}_{\rm ms})$ and a full column rank $\tilde{\mathbf{C}}$. If $n\geq r$, then mean-subtracted DMD is not DFT for Case II in Table 4. If $n\geq r+1$, non-equivalence also holds for Case I.
Proof D.23.
As with Theorem D.20, we prove by contradiction. Assuming we have DMD-DFT equivalence, we can sequentially apply Theorem 4.1 and the identity, $\mathcal{N}(\mathbf{X}_{\rm ms})^{\perp}=\mathcal{R}(\mathbf{X}_{\rm ms}^{H})$, to get the following assertion:
@@
Using Eq. 50, we then get:

### v6

@@ E.2 Equivalence when $\mathbf{X}_{\rm ms}$ has linearly dependent columns
Once again, we can use Eqs. 60 and 59 to introduce $\boldsymbol{\mathbf{l}}^{\perp}$ and leverage its orthogonality to the first $n-1$ columns of $\mathbf{\Theta}_{1}$:
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{\mathbf{C}}_{0}(\mathbf{\Lambda}-\mathbf{I})}\mathbf{\Theta}_{1}\begin{bmatrix}n\\ n-1\\ \vdots\\ 1\end{bmatrix}\left(\frac{-1}{n+1}\right)~{}={}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\boldsymbol{\mathbf{l}}^{\perp})^{H}\mathbf{\Theta}_{1}\begin{bmatrix}n\\ n-1\\ \vdots\\ 1\end{bmatrix}}\left(\frac{-1}{n+1}\right)$
@@
Since we’ve already seen that $\boldsymbol{\mathbf{l}}^{\perp}\neq\boldsymbol{\mathbf{0}}$, the matrix-vector product $\mathbf{X}_{\rm ms}\boldsymbol{\mathbf{e}}_{1}\neq\boldsymbol{\mathbf{0}}$.
Therefore, we have DMD-DFT equivalence.
[…]
E.3 A modal decomposition of mean-subtracted data
When the Koopman mode expansion possesses only a finite number of terms, Lemma B.1 tells us that the time series $\mathbf{Z}$ can be written as the product of $\mathbf{C}$ and $\mathbf{\Theta}$. Here, we derive an analogous factorization for the mean-removed data-set $\mathbf{Z}_{\rm ms}$.
Definition E.6 (A conformal partitioning of $\mathbf{C}$ and $\mathbf{\Theta}$).
Suppose $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ distinct KEFs  and the initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$ is spectrally informative. Additionally, let $1$ be a Koopman eigenvalue i.e., $1\in\{\,\lambda_{i}\,\}_{i=1}^{r}.$ Then, Eq. 36 motivates defining $\boldsymbol{\mathbf{c}}_{1}$, $\mathbf{C}_{\not\owns 1}$ and $\mathbf{\Theta}_{\not\owns 1}$ via the following conformal partitioning:
@@ E.3 A modal decomposition of mean-subtracted data
Table 3: Constructing $\mathbf{C}_{\rm ms}$ and $\mathbf{\Theta}_{\rm ms}$: There are four possible scenarios in mean-subtraction, depending on the presence of a Koopman eigenvalue at 1 and the value of $\boldsymbol{\mathbf{\mu}}$. The peculiarities of each case are incorporated in the definitions of $\mathbf{C}_{\rm ms}$ and $\mathbf{\Theta}_{\rm ms}$ (last two columns), with an eye towards inducing an analogue of Eq. 36 for $\mathbf{Z}_{\rm ms}$.
Case Attributes defining the case Case properties
Spectral property Mean value $\mathbf{C}_{\rm ms}$ $\mathbf{\Theta}_{\rm ms}$
A $1\notin\{\,\lambda_{i}\,\}_{i=1}^{r}$ $\boldsymbol{\mathbf{\mu}}=\boldsymbol{\mathbf{0}}$ $\mathbf{C}$ $\mathbf{\Theta}$
B ′′ $\boldsymbol{\mathbf{\mu}}\neq\boldsymbol{\mathbf{0}}$ $\begin{bmatrix}\mathbf{C}&-\boldsymbol{\mathbf{\mu}}\end{bmatrix}$ $\begin{bmatrix}\mathbf{\Theta}\\ \mathbf{1}_{n+1}^{T}\end{bmatrix}$
C $1\in\{\,\lambda_{i}\,\}_{i=1}^{r}$ $\boldsymbol{\mathbf{\mu}}=\boldsymbol{\mathbf{c}}_{1}$ $\mathbf{C}_{\not\owns 1}$ $\mathbf{\Theta}_{\not\owns 1}$
D ′′ $\boldsymbol{\mathbf{\mu}}\neq\boldsymbol{\mathbf{c}}_{1}$ $\begin{bmatrix}\boldsymbol{\mathbf{c}}_{1}-\boldsymbol{\mathbf{\mu}}&\mathbf{C}_{\not\owns 1}\end{bmatrix}$ $\begin{bmatrix}\boldsymbol{\mathbf{1}}_{n+1}^{T}\\ \mathbf{\Theta}_{\not\owns 1}\end{bmatrix}$
Lemma E.7.
If we construct the two matrices $\mathbf{C}_{\rm ms}$ and $\mathbf{\Theta}_{\rm ms}$ as described by Definition E.6 and Table 3, then, every column of $\mathbf{C}_{\rm ms}$ is non-zero, $\mathbf{\Theta}_{\rm ms}$ has distinct nodes and, more importantly,
@@
Proof E.8.
A direct application of Lemma B.1 to Eq. 11, in the context of Definition E.6 and Table 3.
Remark E.9.
@@
E.4 Over-sampling and linear consistency prevent equivalence
@@
Table 4: Delineating the effect of mean subtraction: There are four possible outcomes of mean-removal, depending on the presence (or absence) of $1$ in the nodes of $\mathbf{\Theta}_{\rm ms}$ and in the Koopman eigen-values. For each case, we detail $\sigma(\mathbf{\Theta}_{\rm ms})$ alongside $\mathbf{C}_{\rm ms}$ and $\mathbf{\Theta}_{\rm ms}$ (whenever they have a simple expression).
Case Attributes defining the case Case properties
Effect of removing $\boldsymbol{\mathbf{\mu}}$ Spectral property $\sigma(\mathbf{\Theta}_{\rm ms})$ $\mathbf{C}_{\rm ms}$ $\mathbf{\Theta}_{\rm ms}$
@@ II $1\in\{\,\lambda_{i}\,\}_{i=1}^{r}$$\{\,\lambda_{i}\,\}_{i=1}^{r}$ $\mathbf{\Theta}$
E.4.1 Auxiliary guarantees
Firstly, we show that a Koopman invariant dictionary ensures the temporal mean, $\boldsymbol{\mathbf{\mu}}$, is zero if and only if we are in Case III.
Lemma E.11.
If $\tilde{\mathbf{C}}$ has full column rank, then, the mean $\boldsymbol{\mathbf{\mu}}$ is zero only in Case III of Table 4.
Proof
The four cases described in Table 4 are mutually exclusive and enumerate all the possible outcomes of mean subtraction. Hence, our objective boils down to establishing that the temporal mean, $\boldsymbol{\mathbf{\mu}}$, is
  1. 1.
    non-zero in Cases I, II and IV.
  2. 2.
    zero in Case III.
We will proceed by contradiction for Cases I and III. The other two cases, II and IV, can be established directly.
Case I
According to Table 4, we have $1\in\sigma(\mathbf{\Theta}_{\rm ms})$ and $1\notin\{\,\lambda_{i}\,\}_{i=1}^{r}$. We will show that the latter condition is contradicted if we have $\boldsymbol{\mathbf{\mu}}=\boldsymbol{\mathbf{0}}$.
Suppose the temporal mean is zero. Then, the additional condition of $1\notin\{\,\lambda_{i}\,\}_{i=1}^{r}$ puts us in Case A of Table 3. Perusing the corresponding value of $\mathbf{\Theta}_{\rm ms}$, we see that it coincides with $\mathbf{\Theta}$ i.e.,
$\boldsymbol{\mathbf{\mu}}~{}=~{}\boldsymbol{\mathbf{0}}\implies\mathbf{\Theta}_{\rm ms}=\mathbf{\Theta}.$
Now, if we draw on the effect of mean-subtraction in Case I of Table 4, we can see that $1$ is a node of $\mathbf{\Theta}_{\rm ms}$ and, hence, also of $\mathbf{\Theta}$:
$\mathbf{\Theta}_{\rm ms}=\mathbf{\Theta}\implies 1\in\sigma(\mathbf{\Theta}).$
However, the simultaneous application of Eqs. 34 and C.1 tells us that the nodes of $\mathbf{\Theta}$ are the Koopman eigenvalues $\{\,\lambda_{i}\,\}_{i=1}^{r}$. Therefore, we have,
$1\in\sigma(\mathbf{\Theta})\implies 1\in\{\,\lambda_{i}\,\}_{i=1}^{r}.$
This contradicts the spectral property of Case I, as listed in Table 4.
Cases II and IV
According to Table 4, the property shared by Cases II and IV is that $1$ is a Koopman eigenvalue i.e., $1\in\{\,\lambda_{i}\,\}_{i=1}^{r}$. We will see that this property, in conjunction with the full column rank of $\tilde{\mathbf{C}}$, ensures a non-zero mean i.e., $\boldsymbol{\mathbf{\mu}}\neq\boldsymbol{\mathbf{0}}.$
To begin with, note that our dictionary lies in the non-redundant span of $r$ distinct Koopman eigenfunctions. So, we can use Eq. 36 to expand the time-average $\boldsymbol{\mathbf{\mu}}$, which is defined in Eq. 10, as follows:
@@ E.4.1 Auxiliary guarantees
Now, $1\in\{\,\lambda_{i}\,\}_{i=1}^{r}$ ensures that $\mathbf{\Theta}\mathbf{1}_{n+1}$ has at least one non-zero coefficient and, hence, is non-zero:
$1\in\{\,\lambda_{i}\,\}_{i=1}^{r}\implies\mathbf{\Theta}\boldsymbol{\mathbf{1}}_{n+1}\neq\boldsymbol{\mathbf{0}}.$
Furthermore, the spectrally informative initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$ lets the full column rank of $\tilde{\mathbf{C}}$ be inherited by $\mathbf{C}$.
$\boldsymbol{\mathbf{\upsilon}}_{1}~{}\textrm{is spectrally informative}~{}\implies\mathbf{C}~{}\textrm{has full column rank.}$.
Combining these two inferences, we get $\boldsymbol{\mathbf{\mu}}\neq\boldsymbol{\mathbf{0}}$:
[…]
Case III
Here, Table 4 tells us that $1\notin\sigma(\mathbf{\Theta}_{\rm ms})$ and $1\notin\{\,\lambda_{i}\,\}_{i=1}^{r}$. We will show that the former is contradicted if the mean is non-zero.
Suppose $\boldsymbol{\mathbf{\mu}}\neq\boldsymbol{\mathbf{0}}$. Taken together with the condition $1\notin\{\,\lambda_{i}\,\}_{i=1}^{r}$, we find ourselves in Case B of Table 3. Perusing the corresponding value of $\mathbf{\Theta}_{\rm ms}$, we see that $1\in\sigma(\mathbf{\Theta}_{\rm ms})$, which is the desired contradiction.
[…]
@@
Proposition E.12.
Suppose $\boldsymbol{\mathbf{\psi}}$ lies in the non-redundant span of $r$ distinct KEFs  and the initial condition $\boldsymbol{\mathbf{\upsilon}}_{1}$ is spectrally informative. Then,
@@
E.4.2 Proof of Theorem 4.5
We are given a dictionary $(\boldsymbol{\mathbf{\psi}})$ that lies in the non-redundant span of $r$ distinct KEFs and a spectrally informative initial condition $(\boldsymbol{\mathbf{\upsilon}}_{1})$. The resulting training data $(\mathbf{X},\mathbf{Y})$ is linearly consistent and over-sampled. Under these conditions, we need to show that mean-subtracted DMD is not a temporal DFT.
We can preface the justification by observing that over-sampling implies well-sampling. So, we can apply Proposition D.3 to deduce that $\tilde{\mathbf{C}}$ has linearly independent columns. Now, we can proceed to a case-by-case analysis.
Case I
Here, Table 4 tells us that $1$ is a node of $\mathbf{\Theta}_{\rm ms}$ i.e., $1\in\sigma(\mathbf{\Theta}_{\rm ms})$. We will establish non-equivalence of $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ and DFT by contradiction.
Suppose mean-subtracted DMD is equivalent to DFT. Then, the vector $\boldsymbol{\mathbf{1}}_{n}^{H}$ lies in the row space of $(\mathbf{\Theta}_{\rm ms})_{1}$. This can be seen by first using Theorem 4.1 to re-cast the equivalence of $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ and DFT as the following identity:
@@
Using Eq. 64, we then get:
[…]
However, over-sampling endows $(\mathbf{\Theta}_{\rm ms})_{1}$ with linearly independent rows. According to Tables 4 and E.9, the matrix $(\mathbf{\Theta}_{\rm ms})_{1}$ has $r+1$ rows and $n$ columns. Since over-sampling means $n\geq r+1$, we can apply Eq. 44 to conclude that $(\mathbf{\Theta}_{\rm ms})_{1}$ has full row rank.
Consequently, $\boldsymbol{\mathbf{d}}^{H}\mathbf{C}_{\rm ms}$ is the unique representation of $\boldsymbol{\mathbf{1}}_{n}^{H}$ in the row space of $(\mathbf{\Theta}_{\rm ms})_{1}$. Taken together with the property that $1$ is a node of $\mathbf{\Theta}_{\rm ms}$ and, hence, of $(\mathbf{\Theta}_{\rm ms})_{1}$, we see that $\boldsymbol{\mathbf{d}}^{H}\mathbf{C}_{\rm ms}$ must be zero at all indices except that corresponding to the node at 1 i.e.,
Now, the sequential application of Lemmas E.11 and E.12 yields a contradiction. Since we are not dealing with Case III, Lemma E.11 gives $\boldsymbol{\mathbf{\mu}}\neq\boldsymbol{\mathbf{0}}$. Therefore, Proposition E.12 says,
$\mathcal{R}(\mathbf{C}_{\rm ms})~{}=~{}\mathcal{R}(\mathbf{C}_{\not\owns 1})~{}=~{}\mathcal{R}\left((\mathbf{C}_{\rm ms})_{\not\owns 1}\right).$
Consequently, we get,
$\boldsymbol{\mathbf{d}}^{H}\mathbf{C}_{\rm ms}~{}=~{}\boldsymbol{\mathbf{0}},$
which contradicts Eq. 65.
Case II
By Table 4, the number $1$ continues to be a node of $\mathbf{\Theta}_{\rm ms}$ i.e, $1\in\sigma(\mathbf{\Theta}_{\rm ms})$.
Non-equivalence of $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ and DFT can be established by re-using the same line of reasoning from Case I, for the most part.
We need only update the arguments used to show $(\mathbf{\Theta}_{\rm ms})_{1}$ has full row rank. Tables 4 and E.9 tell us that $(\mathbf{\Theta}_{\rm ms})_{1}$ has $r$ rows and $n$ columns. Even though $(\mathbf{\Theta}_{\rm ms})_{1}$ has one fewer row than in Case I, over-sampling means $n\geq r+1$, so there are still fewer rows than columns. Consequently, Eq. 44 tells us that $(\mathbf{\Theta}_{\rm ms})_{1}$ has full row rank.
Case III
In this scenario, by Table 4, $1$ is not a node of $\mathbf{\Theta}_{\rm ms}$ i.e., $1\notin\sigma(\mathbf{\Theta}_{\rm ms})$. We can establish the non-equivalence of $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ and DFT by contradiction.
Suppose $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ is equivalent to DFT. Then, the Vandermonde matrix,
$\mathbf{\Theta}_{\rm aug}:=\begin{bmatrix}(\mathbf{\Theta}_{\rm ms})_{1}\\ \mathbf{1}_{n}^{H}\end{bmatrix},$
has linearly dependent rows. To see this, we can retrace the first two steps from Case I - sequentially apply Theorem 4.1 and the identity $\mathcal{N}(\mathbf{X}_{\rm ms})^{\perp}=\mathcal{R}(\mathbf{X}_{\rm ms}^{H})$ to get the following assertion:
[…]
@@
Hence, the rows of $\mathbf{\Theta}_{\rm aug}$ are linearly dependent.
However, this deduction can be contradicted by drawing on the condition of over-sampling. Firstly, $1\notin\sigma(\mathbf{\Theta}_{\rm ms})$ ensures that $\mathbf{\Theta}_{\rm aug}$ has distinct nodes. Now, Table 4 tells us $\mathbf{\Theta}_{\rm aug}$ has $r+1$ rows and $n$ columns. Additionally, the condition of over-sampling ($n\geq r+1$) ensures that there are at-least as many columns as rows. Hence, we can draw on Eq. 44 to infer that $\mathbf{\Theta}_{\rm aug}$ has full row rank which is a contradiction.
Case IV
According to Table 4, the number $1$ continues to be excluded from the nodes of $\mathbf{\Theta}_{\rm ms}$ i.e., $1\notin\sigma(\mathbf{\Theta}_{\rm ms})$.
Non-equivalence of $\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ and DFT can be shown using the same arguments from Case III, with minor updates regarding the size of $\mathbf{\Theta}_{\rm aug}$. Specifically, by Table 4, the matrix $\mathbf{\Theta}_{\rm aug}$ will now have $r$ rows and $n$ columns. Despite having one fewer row than in Case III, over-sampling ($n\geq r+1$) ensures that the columns of $\mathbf{\Theta}_{\rm aug}$ continue to outnumber the rows. Hence, Eq. 44 tells us that $\mathbf{\Theta}_{\rm aug}$ has full row rank.
@@
Remark E.14.
In the above proof of Theorem 4.5, Cases II and IV can also be established under the weaker condition of well-sampling ($n\geq r$). We will leverage this subsequently, to prove Corollary 4.8, in Case II of the just-sampled regime ($n=r$).
E.5 Exploring the under and just-sampled regimes
Proposition E.15.
If the columns of $\tilde{\mathbf{C}}$ are linearly independent and $1$ is a node of $\mathbf{\Theta}_{\rm ms}$, then, $\mathcal{N}(\mathbf{C}_{\rm ms})$ is the one-dimensional subspace spanned by $\mathbf{\Theta}_{\rm ms}\mathbf{1}_{n+1}$.